If John paints a house in 4 hours and Cindy paints a house in 6 hours, how long would it take them if they painted the house together??
HELP!

Answer by Earlsdon(6103)

If John paints the house in 4 hours, then he can paint 1/4 of the house in 1 hour.
If Cindy paints the house in 6 hours, she can paint 1/6 of the house in 1 hour. Together they can paint (1/4 + 1/6 = 5/12) of the house in 1 hour.
So it would take both of them 12/5 hours to paint the entire house.
12/5 hours = 2 hours and 24 minutes.